# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 6.2. Algorithm for Intervals in $\mathbb{N}^{3}$
If $p\in W^{z}\cap X^{z}$, then our treatment of the case $p\in W^{z}$ above yields two relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$, while the our treatment of the case $p\in X^{z}$ yields a third relation $m^{\downarrow}<p$. As $m_{i}^{\downarrow}$, $m_{i+1}^{\downarrow}$, and $m^{\downarrow}$ lie in different components of $\operatorname{dn}(p)$, all three relations can be included in $P$. Since $U$ is an interval in $\mathbb{N}^{3}$, $\operatorname{dn}(p)$ can have at most three components, so no other relations with upper element need to be added to $P$. ∎
Proof of Theorem˜3.17 (i).
We present an algorithm that iterates over $z$ to compute $M_{U_{z}}$, $W^{z}$ and $X^{z}$, as well as the relations $m<p$ of $P$ for $p\in W^{z}\cup X^{z}$.
Inductively, assume that we are given a list $A=M_{U_{z-1}}$, ordered by $x$-coordinate. This will ultimately be updated to $A=M_{U_{z}}$. Let
$B_{z}=\{m\in\mathbb{N}^{2}\mid(m,z)\in M_{U}\}.$
We initialize $X^{z}=W^{z}:=\{\}$ and an empty list $C:=\{\}$ that will contain deleted points from $A$ and is used in the computation of $W^{z}$. For each $m=(m_{x},m_{y})\in B_{z}$ we do the following: Let $L_{m}$ be the prefix of $A$ consisting of points with $y$-coordinate at least $m_{y}$, and let $R_{m}$ be the suffix of $A$ consisting of points with $x$-coordinate at least $m_{x}$. We then add all elements of $\{(n,z)\mid n\in L_{m}\cap R_{m}\}$ to $X^{z}$, along with the associated relations $n^{\downarrow}<(n,z)$. We also store the relation $(m,z)<(n,z)$ for possible inclusion into $P$; see below. We then update
$A\leftarrow(A\setminus(L_{m}\cap R_{m}))\cup\{m\},$
maintaining $A$ in sorted order by $x$-coordinate. The set of points $L_{m}\cap R_{m}$ that are deleted from $A$ are inserted into the list $C$, which is also maintained sorted in $x$-coordinate. Completing the iteration over all elements of $B_{z}$ produces $A=M_{U_{z}}$.
To compute $W^{z}$ and the remaining relations of $P$ whose upper element has third coordinate $z$, we proceed as follows: Writing
$M_{U_{z}}=(m_{1},\ldots,m_{k})$
as above, for each $p=m_{i}\vee m_{i+1}$ where at least one of $m_{i}$ or $m_{i+1}$ is in $M_{U}$, we check whether $p\in W^{z}$. Suppose, as in the definition of $W^{z}$, that $j\in\{i,i+1\}$. Then provided each $m_{j}$ is stored together with the third coordinate of $m_{j}^{\downarrow}$, checking whether $(m_{j},z)\in M_{U}$ is trivial. Checking whether $\nexists\ x\in M_{U_{z-1}}$ with $m_{j}<x<m_{i}\vee m_{i+1}$ amounts to locating $m_{j}$ in the sorted list of points that are deleted from $A$ while computing $M_{U_{z}}$. So, we locate $m_{j}$ in the sorted list $C$ through a binary search and then check if the point in $C$ immediately to the right of $m_{j}$ (the point in $C$ above $m_{j}$ with the least $x$-coordinate if it exists) has the same $y$-coordinate as $m_{j}$. Such a point exists if and only if the second condition in the definition of $W^{z}$ is satisfied for the given choice of $j$. If $m_{i}\vee m_{i+1}\in W^{z}$, then we include the relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$. Once $W^{z}$ has been computed, for each $(n,z)\in X^{z}\setminus W^{z}$, we insert the relation $(m,z)<(n,z)$ computed earlier into $P$.
@@
To explain how we compute $P\cap Q$ from $P$, assume first that $Q$ is specified as input to our algorithm by its upset presentation $(M_{Q},M_{Q^{\prime}})$. We sweep through the 2-D slices $Q^{\prime}_{z}$} of $Q^{\prime}$, computing each $M_{Q^{\prime}_{z}}$ from $M_{Q^{\prime}_{z-1}}$ exactly as we computed $U_{z}$ from $U_{z-1}$ above. For $e=(x,y,z)\in E_{U}$, note that $e\not\in Q$ if and only if $e\in Q^{\prime}$ if and only if $(x,y)\geq m$ for some $m\in M_{Q^{\prime}_{z}}$. The last condition can be checked using at most two binary searches on $M_{Q^{\prime}_{z}}$,
@@
If $Q$ is instead specified by its extrema, a slight variant of the above procedure computes $P\cap Q$ from $P$: Let $U^{\prime}$ denote the upset of the set of maxima of $Q$. We sweep through the 2-D slices $U^{\prime}_{z}$} of $U^{\prime}$, computing each $M_{U^{\prime}_{z}}$ from $M_{U^{\prime}_{z-1}}$. Given $e=(x,y,z)\in E_{U}$ and $M_{U^{\prime}_{z-1}}$, note that $e\in Q$ if and only if $(x,y)\leq m$ for some $m\in M_{U^{\prime}_{z}}$. As above, the second condition can be checked by at most two binary searches on $M_{U^{\prime}_{z}}$,
This completes the specification of our algorithm. We now turn to the complexity analysis. For each $m\in B_{z}$, performing the update of the list $A$ requires time $O(\log|A|)=O(\log|M_{U}|)$; the time cost is dominated by a pair of binary searches in $A$ to find the (possibly null) minimum elements of $L_{m}$ and $R_{m}$. Updating $A$ requires deleting elements of $L_{m}\cap R_{m}$ which can be done in time $O(|L_{m}\cap R_{m}|)$. Thus, this update of $A$ takes $O(\sum_{m\in B}|L_{m}\cap R_{m}|)$. Since each element in $M_{U}$ gets inserted and then possibly deleted only once from $A$ over all levels of $z$, the total update time for $A$ is
$O(\sum_{z}\sum_{m\in B_{z}}|L_{m}\cap R_{m}|)=O(|M_{U}|).$
Similarly, binary searches across all levels of $z$ occur $O(1)$ times for every element in $M_{U}$. Thus, we have $O(|M_{U}|\log|M_{U}|)$ time for all binary searches. Therefore, computing $M_{U_{z}}$ for all values of $z$ takes time
$O(|M_{U}|+|M_{U}|\log|M_{U}|)=O(|M_{U}|\log|M_{U}|)=O(s\log s).$
Computation of $X^{z}$ is dominated by the time to determine the set $L_{m}\cap R_{m}$ for each $m\in B_{z}$. Hence, again this computation takes $O(|M_{U}|\log|M_{U}|)=O(s\log s)$ time over all levels of $z$. Notice that, we check only those points $m_{i}\vee m_{i+1}$ for inclusion in $W^{z}$ where at least one of $m_{i}$ or $m_{i+1}$ is in $M_{U}$. Thus, over all $z$, we check at most $O(|M_{U}|)$ points for inclusion into the set $W^{z}$. Each check for such an inclusion takes only $O(1)$ binary searches in the list $C$. This list can have at most $O(|M_{U}|)$ elements because $C$ contains only elements in $M_{U}$ each of which is deleted from $A$ only once. Thus, computation of $W^{z}$ over all levels of $z$ takes time
$O(|M_{U}|\log|M_{U}|)=O(s\log s).$
@@
To compute $P\cap Q$, we must compute either $M_{Q^{\prime}_{z}}$ or $M_{U^{\prime}_{z}}$ at all values of $z$, which takes time $O(s\log s)$. Besides this, we must perform $O(|E_{U}|)$ binary searches on lists of size $O(s)$. By Theorem˜3.10 (i), we have
$|E_{U}|=O(|M_{U}|)=O(|M_{Q}|)=O(s),$
so the total time required by all of these binary searches is $O(s\log s)$. Thus, the total time to compute $P\cap Q$ from $P$ is $O(s\log s)$. The result follows. ∎
@@ 6.3. Algorithm for Intervals in $\mathbb{N}^{d}$
We next prove Theorem˜3.17 (ii) by giving an algorithm to compute an initial scaffold of an interval in $\mathbb{N}^{d}$ for arbitrary $d$, and bounding its complexity. As in the $d=3$ case, our strategy is to first compute an initial scaffold $P$ of $U=\operatorname{up}[Q,\mathbb{N}^{d}]$ and then compute the intersection $P\cap Q$, which is a hull skeleton of $Q$ by Proposition˜6.1.
To compute $P$, we use following structural result:
@@
For $T\subseteq U$ the full subposet consisting of elements of $M_{U}$ and their pairwise joins, $P\subset U$ is an initial scaffold of $U$ if and only if it is an initial scaffold of $T$.
@@
It is easily checked that any element of $E_{U}$ is a pairwise join of elements in $M_{U}$. (Equivalently, any element in the support of $\beta_{1}^{\hat{U}}$ is such a join, which is immediate by considering the Taylor resolution of $\hat{U}$ [miller2005combinatorial].) Hence $E_{U}\subseteq T$. We claim that for any $t\in T$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ induces a bijection on path components. The claim implies that $I_{T}=I_{U}$. Since $M_{T}=M_{U}$, the claim then also implies that $T$ and $U$ have the same initial scaffolds, as desired.
@@
It remains to prove the claim. Since $M_{T}=M_{U}$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ clearly induces a surjection on path components. We need to show that the induced map is also an injection. Suppose $w,y\in T$, with $w$ and $y$ contained in the same component of $\operatorname{dn}(t,U)$. It suffices to show that $w$ and $y$ are contained in the same component of $\operatorname{dn}(t,T)$. Our argument is reminiscent of the proof of Lemma˜5.11. By assumption, there exists a path $w=q_{1},\ldots,q_{n}=y$ from $w$ to $y$ in $\operatorname{dn}(t,U)$. We may assume without loss of generality that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either $q_{i-1}\leq q_{i}\leq q_{i+1}$ or $q_{i-1}\geq q_{i}\leq q_{i+1}$. In the former case, we call $q_{i}$ an upper point, and in the latter case, we call $q_{i}$ a lower point. Moreover, we may assume without loss of generality that $n>2$, and that $q_{2}$ and $q_{n-1}$ are lower points. Replacing each lower point with a minimum below it, and then replacing each upper point $q_{i}$ with the join $q_{i-1}\vee q_{i+1}$, we obtain a path from $w$ to $y$ in $\operatorname{dn}(t,T)$. Thus, $w$ and $y$ lie in the same component of $\operatorname{dn}(t,T)$, as we wanted. ∎
Proof of Theorem˜3.17 (ii).
@@
By Proposition˜6.5, to compute an initial scaffold of $U=\operatorname{up}[Q]$, it suffices to compute an initial scaffold of $T$. To do this, we first compute $T\setminus M_{T}$ and then order this set in a way compatible with the partial order induced by $\mathbb{N}^{d}$, e.g., lexicographically. We then iterate through the elements of $T\setminus M_{T}$ in order, adding each $t\in E_{T}$ to $P$, together with the required relations with upper element $t$. To elaborate, let $P^{t}$ be the portion of $P$ computed so far when we come to $t\in T\setminus M_{T}$. We compute $P^{t}\cap\operatorname{dn}(t,T)$ by comparing each element of $P^{t}$ to $t$. Similar to our algorithm for computing initial scaffolds of general finite posets given in Theorem˜3.13, we then compute the components of $P^{t}\cap\operatorname{dn}(t,T)$ via depth-first search. Proposition˜4.7 implies that it suffices here to consider the components of $P^{t}\cap\operatorname{dn}(t,T)$ rather than the components of $\operatorname{dn}(t,T)$. If $P^{t}\cap\operatorname{dn}(t,T)$ has more than one component, then we add $t$ to $P$, and for each component $A$ of $P^{t}\cap\operatorname{dn}(t,T)$, we chose a minimum $m\in A$ and add the relation $m<t$ to $P$.
@@
For each element $t\in T\setminus M_{T}$ considered in the above algorithm, identifying the components of $P^{t}$ requires time $O(|P|)=O(|M_{Q}|^{2})=O(s^{2})$. At most $d$ edges are added to $P^{t}$ per element $t$, so adding these edges requires constant time. Since $T$ has size $O(|M_{Q}|^{2})=O(s^{2})$, the total time to compute an initial scaffold of $P$ is $O(s^{4})$.
@@
Given $P$, we can straightforwardly compute $P\cap Q$ in time $O(s^{3})$. To explain, if the input to our algorithm is the upset presentation $(M_{Q},M_{Q^{\prime}})$, then for each $e\in E_{U}$, checking whether $e\in P$ amounts to checking whether $e\geq m$ for some $m\in M_{Q^{\prime}}$, as we have seen in the proof of Theorem˜3.17 (i). Since $|M_{Q^{\prime}}|=O(s)$ and $|E_{U}|=O(s^{2})$ by Theorem˜3.10 (i), the total time required for all such comparisons is $O(s\cdot s^{2})=O(s^{3})$. If $Q$ is instead given by the minima and the maxima of $Q$, the computation of $P\cap Q$ is analogous. Thus, the total time to compute an initial scaffold of $Q$ is $O(s^{4})$. ∎

### v2

@@ 6.2. Algorithm for Intervals in $\mathbb{N}^{3}$
If $p\in W^{z}\cap X^{z}$, then the case $p\in W^{z}$ treated above yields two relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$, while the case $p\in X^{z}$ yields a third relation $m^{\downarrow}<p$. As $p\in X^{z}$, there is no $l<m$ with $(l,z-1)\in\operatorname{dn}(p)$, so $m_{i}^{\downarrow}$ and $m_{i+1}^{\downarrow}$ both have third coordinate $z$. Thus, $m_{i}^{\downarrow}$, $m_{i+1}^{\downarrow}$, and $m^{\downarrow}$ lie in different components of $\operatorname{dn}(p)$, hence all three relations can be included in $P$. Since $\operatorname{dn}(p)$ can have at most three components, no other relations with upper point $p$ need to be added to $P$. ∎
To prepare for our algorithm, we recall that a self-balancing binary search tree $\mathcal{T}$, such as a red-black tree [14, Chapter 13], stores an ordered set of size $l$ in a way that supports binary searches, insertions, and deletions, each in $O(\log l)$ time. In addition, $\mathcal{T}$ supports range queries: Given keys $a$ and $b$, it returns the set $S\coloneqq\{x\in\mathcal{T}\mid a\leq x\leq b\}$ as an ordered list in time $O(\log l+|S|)$; see [14, Exercise 12.2-8].
Proof of Theorem˜3.16 (i).
Given $M_{U}=M_{Q}$, our algorithm first computes an initial scaffold $P$ of $U$ as in Proposition˜6.3 by sweeping through the planar slices $U_{z}$ of $U$. Initially, we set $P\leftarrow M_{U}$, with no non-identity relations. As the algorithm proceeds, it adds points and relations to $P$.
Let
$T^{z}\coloneqq\{m^{\downarrow}\mid m\in M_{U_{z}}\},$
and note that $M_{U_{z}}$ is the image of $T^{z}$ under coordinate projection onto the xyx-y plane. We maintain $T^{z}$, ordered by $x$-coordinate, using a self-balancing binary search tree $\mathcal{T}$. As we modify $\mathcal{T}$ during the course of our algorithm, $\mathcal{T}$ will always maintain the property that its image under coordinate projection onto the xyx-y plane is pairwise incomparable. Thus, the order of $\mathcal{T}$ by $x$-coordinate is the same as the opposite of the order by $y$-coordinate.
Inductively, assume that $\mathcal{T}=T^{z-1}$ has been computed. Let
$Y^{z}=\{m\in\mathbb{N}^{2}\mid(m,z)\in M_{U}\}.$
We store $Y^{z}$ in a list ordered by $x$-coordinate. Using $T^{z-1}$ and $Y^{z}$, we compute $M_{U_{z}}$, $W^{z}$, $X^{z}$, the $X^{z}$-relations, and the $W^{z}$-relations.
First, we compute $X^{z}$ and a set of relations of $P$ which includes all of the $X^{z}$-relations, as well as some of the $W^{z}$-relations. To do so, we initialize an empty self-balancing binary search tree that will store the points of $X^{z}$, ordered by $x$-coordinate. We iterate through the points of $Y^{z}$ in any order. For each $m=(m_{1},m_{2})\in Y^{z}$ we do the following: Let $L$ be the prefix of $\mathcal{T}$ consisting of points with $y$-coordinate at least $m_{2}$, and let $R$ be the suffix of $\mathcal{T}$ consisting of points with $x$-coordinate at least $m_{1}$. We determine the minimum point $a$ of $R$ (null if $R$ is empty) and maximum point $b$ of $L$ (null if $L$ is empty) by a pair of binary searches on $\mathcal{T}$. Given $a$ and $b$, we can compute $L\cap R$ by a range search on $\mathcal{T}$. We then add each point of
$\{(m^{\prime},z)\mid(m^{\prime},t)\in L\cap R\}$
to $X^{z}$ (hence to $P$), and also add both of the relations $(m^{\prime})^{\downarrow}<(m^{\prime},z)$ and $(m,z)<(m^{\prime},z)$ to $P$. (Note that the former relation is an $X^{z}$-relation. If $m^{\prime}$ and $m$ share no coordinate, then the latter relation is also an $X^{z}$-relation; otherwise, it is either an $X^{z}$-relation or a $W^{z}$-relation.) Using insertion and deletion operations, we then set
$\mathcal{T}\leftarrow(\mathcal{T}\setminus(L\cap R))\cup\{(m,z)\}.$
After iterating over all points of $Y^{z}$, we have $\mathcal{T}=T^{z}$, and moreover, the computation of $X^{z}$ and all $X^{z}$-relations is complete.
To compute $W^{z}$ and the remaining $W^{z}$-relations, we proceed as follows: As above, we write
$M_{U_{z}}=(m_{1},\ldots,m_{k}).$
We initialize an integer variable $\textsc{SkipIndex}=0$, which we use to avoid processing the same element of $W^{z}$ twice. Noting that $Y^{z}\subseteq M_{U_{z}}$, for each $m_{i}\in Y^{z}$ in increasing order, we do the following:
  • If $i<k$, then we check via binary search whether there exists $(v,z)\in X^{z}$ with $v_{2}=(m_{i})_{2}$ and $v_{1}<(m_{i+1})_{1}$. It is easy to show that
    $p\coloneqq(m_{i}\vee m_{i+1},z)\in W^{z}$
    if and only if no such $(v,z)$ exists. Thus, if $(v,z)$ does not exist, then we do the following:
    • If $p\not\in X^{z}$, then we add $p$ to $P$, along with the relations $m^{\downarrow}_{i}<p$ and $m_{i+1}^{\downarrow}<p$. If $p\in X^{z}$, then one of these two relations is already in $P$; we add the other to $P$.
    • We set $\textsc{SkipIndex}\leftarrow i+1$.
  • Symmetrically, if $i>1$ and $\textsc{SkipIndex}\neq i$, then we check whether there exists $(v,z)\in X^{z}$ with $v_{1}=(m_{i})_{1}$ and $v_{2}<(m_{i-1})_{2}$. We have
    $p\coloneqq(m_{i-1}\vee m_{i},z)\in W^{z}$
    if and only if no such $(v,z)$ exists. Thus, if $(v,z)$ does not exist, then we do the following:
    • If $p\not\in X^{z}$, then we add $p$ to $P$, along with the relations $m^{\downarrow}_{i-1}<p$ and $m_{i}^{\downarrow}<p$. If $p\in X^{z}$, then one of these two relations is already present in $P$; we add the other to $P$.
This completes the computation of $P$.
@@
To explain how we compute $P\cap Q$ from $P$, assume first that $Q$ is specified as input to our algorithm by its upset presentation $(M_{Q},M_{Q^{\prime}})$. We sweep through the planar slices $Q^{\prime}_{z}$} of $Q^{\prime}$, computing each $M_{Q^{\prime}_{z}}$ from $M_{Q^{\prime}_{z-1}}$ exactly as we computed $M_{U_{z}}$ from $M_{U_{z-1}}$ above. For $e=(x,y,z)\in E_{U}$, note that $e\not\in Q$ if and only if $e\in Q^{\prime}$, and that $e\in Q^{\prime}$ if and only if $(x,y)\geq m$ for some $m\in M_{Q^{\prime}_{z}}$. The last condition can be checked using at most two binary searches on $M_{Q^{\prime}_{z}}$.
@@
If $Q$ is instead specified by its extrema, a slight variant of the above procedure computes $P\cap Q$ from $P$: Let $U^{\prime}$ denote the upset of the set of maxima of $Q$. We sweep through the planar slices $U^{\prime}_{z}$} of $U^{\prime}$, computing each $M_{U^{\prime}_{z}}$ from $M_{U^{\prime}_{z-1}}$. For $e=(x,y,z)\in E_{U}$, note that $e\in Q$ if and only if $(x,y)\leq m$ for some $m\in M_{U^{\prime}_{z}}$. As above, the second condition can be checked by at most two binary searches on $M_{U^{\prime}_{z}}$.
This completes the specification of our algorithm. Its correctness follows straightforwardly from Proposition˜6.3.
We now turn to the complexity analysis. We first consider the cost of computing $P$. Note that for each $m\in M_{U}=M_{Q}$, we have $m\in Y^{z}$ for exactly one value of $z$. For each $m\in Y^{z}$, we perform $O(1)$ binary searches on $\mathcal{T}$ and $O(1)$ binary searches on $X^{z}$. Since $|\mathcal{T}|\leq|M_{Q}|$, $|X^{z}|\leq|M_{Q}|$, and $|\{z\in\mathbb{N}\mid Y^{z}\neq\emptyset\}|<|M_{Q}|$}, the total time required by these binary searches across all levels $z$ is $O(|M_{Q}|\log|M_{Q}|)$. In addition, across all levels $z$, each element of $M_{Q}$ is inserted into $\mathcal{T}$ exactly once and removed from $\mathcal{T}$ at most once. Thus, the total cost of all such insertions and deletions is again $O(|M_{Q}|\log|M_{Q}|)$. For each $m\in Y^{z}$, the associated range search takes time $O(\log|\mathcal{T}|+s_{z})=O(\log|M_{Q}|+s_{z})$, where $s_{z}$ is the number of elements removed from $\mathcal{T}$ at level $z$. Since $\sum_{z}s_{z}=O(|M_{Q}|)$, the total cost of all range searches is $O(|M_{Q}|\log|M_{Q}|).$ Beside these operations, computing $P$ requires $O(1)$ time per element of $Y^{z}$. Hence the total time to compute $P$ is $O(|M_{Q}|\log|M_{Q}|)=O(n\log n)$.
@@
To compute $P\cap Q$, we must compute either $M_{Q^{\prime}_{z}}$ or $M_{U^{\prime}_{z}}$ at all values of $z$, which takes time $O(n\log n)$. Besides this, we must perform $O(|E_{U}|)$ binary searches on lists of size $O(n)$. By Theorem˜3.10 (i), we have $|E_{U}|=O(|M_{Q}|)=O(n),$ so the total time required by all of these binary searches is $O(n\log n)$. Thus, the total time to compute $P\cap Q$ from $P$ is $O(n\log n)$. The result follows. ∎
Remark 6.5.
In view of Lemma˜5.12, the algorithm underlying our proof of Theorem˜3.16 (i) specializes to an $O(n\log n)$-time algorithm for computing the support of the first Betti numbers of a monomial ideal $J\subset R^{3}$ with $n$ minima. A slight modification of the algorithm yields the first Betti numbers themselves. We expect that the incremental approach of our algorithm also extends to compute a minimal resolution of $J$ (and in particular, the second Betti numbers of $J$), in the same asymptotic time. An outline of a rather different algorithm for this has previously been given in [35, Section 3], without explicit complexity bounds.
@@ 6.3. Algorithm for Intervals in $\mathbb{N}^{d}$
We next prove Theorem˜3.16 (ii) by giving an algorithm to compute an initial scaffold of an interval in $\mathbb{N}^{d}$, for arbitrary $d$. As in the $d=3$ case, our strategy is to first compute an initial scaffold $P$ of $U=\operatorname{up}[Q,\mathbb{N}^{d}]$ and then compute $P\cap Q$, which is an initial scaffold of $Q$ by Proposition˜6.1.
To compute $P$, we use the following structural result:
@@
For $T\subseteq U$ the full subposet consisting of points of $M_{U}$ and their pairwise joins, $P\subseteq U$ is an initial scaffold of $U$ if and only if it is an initial scaffold of $T$.
@@
It is easily checked that any point of $E_{U}$ is a pairwise join of points in $M_{U}$. (Equivalently, any point in the support of $\beta_{1}^{\hat{U}}$ is such a join, which is immediate by considering the Taylor resolution of $\hat{U}$ [35].) Hence $E_{U}\subseteq T$. We claim that for any $t\in T$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ induces a bijection on components. The claim implies that $E_{U}=E_{T}$, and hence that $I_{U}=I_{T}$. Since $M_{U}=M_{T}$, it then follows that $U$ and $T$ have the same initial scaffolds, as desired.
@@
It remains to prove the claim. Our argument is quite similar to the proof of Lemma˜5.11. Since $U$ has finite downsets, every component of $\operatorname{dn}(t,U)$ contains a minimum. Since $M_{T}=M_{U}$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ therefore induces a surjection on components. We need to show that the induced map is also an injection. Suppose $w,y\in T$, with $w$ and $y$ contained in the same component of $\operatorname{dn}(t,U)$. It suffices to show that $w$ and $y$ are contained in the same component of $\operatorname{dn}(t,T)$. By assumption, there exists a path $w=q_{1},\ldots,q_{k}=y$ from $w$ to $y$ in $\operatorname{dn}(t,U)$. We may assume without loss of generality that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either
$q_{i-1}\leq q_{i}\geq q_{i+1}\quad\textup{or}\quad q_{i-1}\geq q_{i}\leq q_{i+1}.$
We call $q_{i}$ an upper point in the former case and a lower point in the latter case. Moreover, we may assume without loss of generality that $k>2$, and that $q_{2}$ and $q_{k-1}$ are lower points. Replacing each lower point that is not a minimum with a minimum below it, and then replacing each upper point $q_{i}$ with the join $q_{i-1}\vee q_{i+1}$, we obtain a path from $w$ to $y$ in $\operatorname{dn}(t,T)$. Thus, $w$ and $y$ lie in the same component of $\operatorname{dn}(t,T)$, as we wanted. ∎
Proof of Theorem˜3.16 (ii).
By Proposition˜6.6, to compute an initial scaffold $P$ of $U=\operatorname{up}[Q]$, it suffices to compute an initial scaffold of $T$. To do this, we initialize $P\leftarrow M_{U}$, with no non-identity relations. We then compute $T\setminus M_{T}$, ordered in a way compatible with the partial order induced by $\mathbb{N}^{d}$, e.g., lexicographically.
@@
We next iterate through the points of $T\setminus M_{T}$ in order, adding each $t\in E_{T}$ to $P$, together with the required relations with upper point $t$. To elaborate, let $P^{t}$ be the part of $P$ computed so far when we come to $t\in T\setminus M_{T}$. By Proposition˜4.8, to decide whether $t\in E_{T}$ and, if so, to compute the relations of $P$ with upper point $t$, it suffices to compute the components of ${P^{t}\cap\operatorname{dn}(t,T)}=P\cap\operatorname{dn}(t,T)$ rather than those of $\operatorname{dn}(t,T)$. We compute $P^{t}\cap\operatorname{dn}(t,T)$ by comparing each point of $P^{t}$ to $t$. Similarly to our algorithm for computing initial scaffolds of general finite posets given in the proof of Theorem˜3.13, we then compute the components of $P^{t}\cap\operatorname{dn}(t,T)$ via depth-first search. If $P^{t}\cap\operatorname{dn}(t,T)$ has more than one component, then we add $t$ to $P$, and for each component $A$ of $P^{t}\cap\operatorname{dn}(t,T)$, we choose a minimum $m\in A$ and add the relation $m<t$ to $P$.
@@
For each point $t\in T\setminus M_{T}$ considered in the above algorithm, identifying the components of $P^{t}$ requires time $O(|P|)=O(|M_{Q}|^{2})=O(n^{2})$, where the first equality follows from Theorem˜3.10 (ii). At most $d$ relations are added to $P^{t}$ per point $t$, so adding these relations requires constant time. Since $T$ has size $O(|M_{Q}|^{2})=O(n^{2})$, the total time to compute the initial scaffold $P$ is $O(n^{4})$.
@@
Given $P$, we can straightforwardly compute $P\cap Q$ in time $O(n^{3})$. To explain, if the input to our algorithm is the upset presentation $(M_{Q},M_{Q^{\prime}})$, then for each $e\in E_{U}$, checking whether $e\in Q$ amounts to checking whether $e\geq m$ for some $m\in M_{Q^{\prime}}$, as in the proof of Theorem˜3.16 (i). Since $|M_{Q^{\prime}}|=O(n)$ and $|E_{U}|=O(n^{2})$ by Theorem˜3.10 (ii), the total time required for all such comparisons is $O(n\cdot n^{2})=O(n^{3})$. If $Q$ is instead given by the extrema of $Q$, then the computation of $P\cap Q$ is analogous. Thus, the total time to compute an initial scaffold of $Q$ is $O(n^{4})$. ∎